OpenStudy (anonymous):

Solve the system of equations. 5x - y = 3 and -10x + 2y = -6

OpenStudy (y2o2):

no solution

OpenStudy (msmr):

actually, wouldn't it be any real number that satisfies one of the systems?

OpenStudy (msmr):

because if y=5x-3 (from the first equation) and you substitute it into the next equation it becomes: -10x+2(5x-3)=-6 -10x+10x-6=-6 0=0

OpenStudy (msmr):

if x is three, then y =15-3=12 plug that in to the other equation and: -10x+24=-6 -10x=-24 x=2.4 so you can pick any number for one of the variables and substitute it in for the other one